OpenStudy (anonymous):

Write the standard form of the line that passes through the given points. (6, 1) and (5, 4)

OpenStudy (anonymous):

@Love_Ranaa

OpenStudy (thatgingeronfire):

Alright so first things first, let's find the slope. Slope of a line is \[m=y _{2}-y _{1}/x _{2}-x _{1}\] 6 and 5 are x's and 1 and 4 are y's. We need to keep them in order, so we'll do \[m=4-1/5-6\] \[m=-3/1=-3\] So the slope is -3

OpenStudy (anonymous):

But I need the standard form

OpenStudy (thatgingeronfire):

I know I'm getting to that

OpenStudy (thatgingeronfire):

So find the y-intercept with the slope. We have y=-3x+b use a point to plug in for the x and y \[(1)=-3(6)+b\] -3*6 is -18, so we have 1=-18+b +18 to both sides, and we have 19=b The y-intercept is (0,19) \[y=-3x+19\] Now we have the slope-intercept form, we can turn it into the standard form.

OpenStudy (thatgingeronfire):

So the standard form is \[Ax+By=C\] we want to isolate the c, so we'll do +3x to both sides. \[y+3x=-3x+3x+19\] This now becomes \[y+3x=19\] or \[3x+y=19\] So do you understand how to get it?

OpenStudy (anonymous):

No lol

OpenStudy (thatgingeronfire):

What do you not get